Prev Next

SW&IT Meeting Minutes 2021-03-29

Attendees: David, Matthieu, Michael, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20210322
LR20210322 - Bec, Matthieu 40.0
LR20210322 - Edmeades, David 42.0
LR20210322 - Sizemore, Michael 0.0
LR20210322 - Grenz, Paul 0.0
LR20210322 - Hooper, Stephen 80.0
LR20210322 - Zhang, Yang 40.0
LR20210322 - Zhang, Xianyu 40.0

News from management

  • 2 candidates moving to tech interviews

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • IT 8344 (LUCI1 RTD on Obs2 error, and bitIndex). Barry commented (in instrument meeting) that Aladdin for Luci is a special version. If we want source code, we may need to ask other people (he referenced Alexander).
  • IT 8353 (Earch orientation data). Daily version seems to be only available from NASA (so far). Should we go for it? And if so, we need to register (better to create an offial account "tcs@lbto.org" than use personal email and registration).

Team reports

MB

  • Last week
    • OC, Management
    • ATT PM FTE & priorities
    • OT demo for SHARK NIR, went out for evaluation purpose
  • This week
    • ATT PM FTE & priorities
    • candidates tech questionaire and follow up interviews

YZ

  • Previous:
    • Released a new version of RB_Science program.
    • Some TMS work (utility update in progress, and discussion during meeting).
    • Continuing some dms osco work.
    • We finally had a new OSCO dataset (and the live page work as expected).
    • Cravat/OSCO meeting/wiki work after some pause (and some new work).
    • Some work on IT 8353 (PCS earch orientation data problem).
  • Next:
    • Hopefully finish dms osco fix and tweaking, and then make a public announcement.
    • Some OSCO background process upgrade (and if have time, look into things discussed during the meeting).
    • Release TMS utility (and wiki work). Help Heejoo analying recent data.
    • Drafting TMS PC software redesign requests.
    • Start TMS database transition.
    • Look into MODS 64-bit stuff more.
    • Safety training.

XZ

  • Last week
    • ATT AdSec/SW tasks
    • plan for DX ASM
    • DSPCoil Current issue
    • AO support
  • This week
    • ATT AdSec /SW
    • Safety training
    • AO support

DE

  • Syno support prep
  • Sunday office trip for Synology, no work done because of communication failures internal to Synology support
  • Much back-and-forth about asset management, many incorrect entries in the records
  • Saftey training

Next:
  • Revert support access changes
  • Server room HVAC fan too high; will contact FM
  • Cisco
  • Resource usage spreadsheet

PSG

  • Last week
    • Gathering information about D425 controller
    • Finding flash card /backup for D425.
    • Tested UMAC software for AGW
  • This week
    • Fixing AGW UMAC code, removing TCS dependencies
    • Testing UMAC code again
    • Downloading D425 config from spare unit as test?

SH

  • DIMM with LeRoy
  • UITS stuff
  • Some UA Secure
  • IT meeting
  • Dave T.
  • Slack and e-mail groups with Laura.
  • core-a problems.
  • Vet and half-day Friday.

next:

  • PEPSI VM
  • some more VM stuff
  • DIMM with LeRoy (maybe)
  • LDAP
Topic revision: r9 - 05 Apr 2021, StephenHooper
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback